OpenStudy (anonymous):

perform the indicated operation (8+2i)(8-2i)

OpenStudy (anonymous):

8*8-8*2i+2i*8-2i*2i 64+4 68

OpenStudy (anonymous):

i*i=-1

OpenStudy (anonymous):

(a+bi)(a-bi)=a^2+b^2 the product of a complex number and its conjugate is a sum of squares; memorize this formula